Controls Meeting 3-Apr-2014
- Minutes from previous meeting Controls Meeting 20-Mar-2014
- BCAL pulser control (Orlando)
- Alarms for voltages (Nerses)
- media:BCAL_Interlock.pptx | Status of BCAL interlocks (Dave)
- Radiator insertion stage test (Hovanes, Yi)
- Status of the cabling of the motorized stages (Scot)
- Status of FDC/CDC Gas System (Scot,Dave)
- Any other business
9:30AM Thurs 3-Apr-2014 CC A110
Present: Dave, Mark S., Scot, Orlando, Hovanes, Elton, Nerses
- Orlando showed the status of the BCAL pulser control.
- The application is pretty much ready to be used.
Alarms for voltages
- Nerses is working on the scheme to generate alarms for voltages in EPICS.
- He showed the logic he is currently considering.
- There was a suggestion to have an alarm when the voltages are set to be off, but they are still on,
- Orlando pointed out that there will be a delay between the moment the setpoint is set and setpoint is readback. This time delay needs to be taken into account.
- Dave reported significant progress on the BCAL interlocks .
- Upstream is wired and tested. Downstream interlock is in progress.
- Downstream chiller is not connected as of today. Will be connected most likely by the end of today.
- Buttons need to be change to match Elton's convention .
- Hovanes asked Elton if the expert GUI that Dave has will need to be copied to EPICS or EPICS will need to have only a GUI for the shift personnel with very limited personality. This was followed by a discussion about the need to have a PLC in the counting house.
- The test were done by Yi and Hovanes. The stage satisfied the specs. The limit switches will have to replaced.
- Slava and Hovanes handed the stage to Tim for installation on the tagger hall. The PS converter stage is in the FEL now.
Cabling of the motors
- Tina has been working on the cables for the Hall D motorized stages.
- Collimator cable will be complete today.
- Scot needs the notation for the low and high limit switches for each stage.
Gas system controls
- Gas system is pretty close to be complete.
- Procedures are being circulated for the controls.
- Dave reported that 80% of the PLC GUIs are complete.
- Michael Staib from CMU is working on the EPICS controls screens for the gas system .
- Elton would like to purchase flow meters with remote readout. He will send an e-mail with the data-sheet or the specification for the flow-meter he liked.
- Hovanes will send info to Scot on limit switches for motors.
- Beni will forward Hovanes the draft of the procedures for the gas system to Hovanes.
- Dave will change the buttons on the BCAL temperature GUIs to match Hall D convention for controls buttons.
- Elton will send an e-mail with the information on the flow meter.
- Dave will check if the flow-meter that Elton wants could be easily integrated into the PLC controls.
- Orlando will commit the final version of the BCAL pulser control and Hovanes will check out the GUIs on the official tree.